# SlideChat: A Large Vision-Language Assistant for Whole-Slide Pathology Image Understanding

### v2

@@ 3.3 Two-stage Training
SlideChat adopts a two-stage training approach (see Figure 2 B). In the first stage, the primary objective is to align the large language model’s (LLM) word embeddings with the visual features extracted from whole slide images. This alignment enables the LLM to interpret visual representations from the slide-level encoder, facilitating the effective utilization of the intricate features within the slides. During this stage, SlideChat is trained to generate descriptive captions using 4.2K WSI-caption pairs from SlideInstruction. Specifically, only the slide-level encoder and projection matrix are updated, while the patch-level encoder and LLM weights remain fixed.
@@
In the second stage, we focus on visual question-answering tasks to train the model to accurately respond to domain-specific questions concerning whole slide images. During this phase, the model develops the ability to handle a broad range of multimodal instructions, enabling it to generate answers by effectively integrating both visual and textual information. For example, the model must perform various pathology tasks, such as describing the extent of tumor invasion or assessing the degree of cellular differentiation. To accomplish this, we utilize 176K WSI VQAs from SlideInstruction in the second training stage, allowing the slide encoder, projection layer, and large language model components to be fully trainable to ensure comprehensive adaptability. This training approach significantly enhances the model’s capability to handle diverse pathology-related tasks, thereby increasing its effectiveness in real-world clinical and research settings.
Table 2: Captioning performance across different methods on SlideBench-Caption. Slide (T) refers to the WSI thumbnail with size of $1024\times 10241024 × 1024$.
MLLMs Input BLEU-1 BLEU-2 BLEU-3 BLEU-4 Rouge-L GPT-score
@@
4 Experiment
We conducted following experiments to evaluate three key aspects of SlideChat: (1) its whole slide image captioning capability, which assesses proficiency in generating descriptive captions that accurately summarize the critical pathological features of a WSI; (2) its visual question-answering (VQA) ability across various complex pathological scenarios and its generalizability in zero-shot settings; and (3) SlideChat’s ability to process gigapixel WSIs, capturing both essential global context and intricate details, thereby enhancing its performance compared to patch-level multimodal large language models. For WSI captioning baselines, we benchmark against MI-Gen (Chen et al., 2023), a state-of-the-art method specifically designed for this task. Given that existing MLLMs cannot handle the gigapixel scale of whole slide images, we establish baseline comparisons using two approaches: (1) randomly selecting 30 patches from each WSI and inputting them into MLLMs (e.g., GPT-4 (Achiam et al., 2023), LLaVA-Med (Li et al., 2024a), MedDr (Li et al., 2024a)), followed by a majority voting scheme to generate slide-level predictions; and (2) directly inputting a WSI thumbnail, resized to 1024×1024 pixels, into the MLLMs. For VQA tasks, we further evaluate performance by comparing against random prediction baselines and text-only models, thereby assessing the incremental contribution of visual information. Unless otherwise specified, SlideChat is configured with the patch-level encoder CONCH (Lu et al., 2024a), the slide-level encoder LongNet (Ding et al., 2023), and utilizes the Qwen2.5-7B-Instruct (Yang et al., 2024) as LLM.
SlideBench-Caption
We report BLEU, ROUGE, and GPT scores to evaluate caption generation performance in Table 2. For the GPT score, we use GPT-4 to assess the similarity between the generated captions and the ground truth, providing an overall score on a scale of 1 to 10, with higher scores indicating better performance. When utilizing patch-level inputs, GPT-4o generates individual descriptions for each patch, which are subsequently integrated to create the final slide-level caption. However, this approach yields poor performance, as evidenced by a BLEU-1 score of 0.16 and a GPT-score of 1.54. These results suggest that the patch-based method fails to adequately capture the broader context necessary for accurate WSI captioning. When the WSI thumbnail of size 1024×1024 pixels is used as input to GPT-4o, performance decreases further, with a BLEU-1 score of 0.10 and a GPT-score of 1.03. This suggests that while the thumbnail offers a global view of the slide, it may lack the resolution and detail necessary for generating precise and informative captions. In contrast, MI-Gen, a model specifically designed for WSI captioning, demonstrates significantly superior performance across all metrics, achieving a BLEU-1 score of 0.37 and a GPT score of 4.14. Similarly, SlideChat, shows competitive results with a BLEU-1 score of 0.37 and a GPT score of 4.11. These outcomes highlight SlideChat’s ability to effectively integrate both local and global information from the slides and confirm its efficiency in describing whole-slide images, as illustrated by several examples shown in Figure 7 in the Appendix.
SlideBench-VQA (TCGA)
We further evaluate SlideChat’s overall performance on the multiple-choice VQA benchmark. The results on SlideBench-VQA (TCGA), presented in Table 3, compare different methods across three domains: Microscopy, Diagnosis, and Clinical, along with an overall performance score. Random selection achieves an overall score of 25.02% accuracy, serving as a baseline for answer distribution but demonstrating poor performance. While GPT-4, relying solely on text input, outperforms random predictions, it continues to struggle with accurately answering questions. When GPT-4o incorporates patch-level inputs, its performance improves markedly, reaching a score of 57.91% and underscoring the crucial role of detailed visual data. However, using a WSI thumbnail results in a lower score of 34.07%, as the reduced detail restricts its ability to deliver precise answers. MedDr performs well, achieving an overall score of 67.70% with patch-level inputs, though this drops slightly to 64.25% when using the slide thumbnail due to the loss of fine visual details. SlideChat outperforms all other methods, attaining a leading overall accuracy of 81.17%, excelling across all categories and significantly surpassing the competition. Even in more fine-grained pathological scenarios, as depicted in the left portion of Figure 4, SlideChat remains the top-performing model across 13 tasks, particularly in areas such as cytomorphological characteristics, histopathological changes, disease detection, disease classification, and staging and grading, which require the identification of complex pathological visual features. Compared with baselines taking some patches or slide thumbnial as inputs, SlideChat has the capability to analyze a significantly greater number of pathological features with enhanced detail, effectively capturing both localized features and overarching global patterns, allowing SlideChat to provide more accurate and nuanced insights into pathological variations. In Figure 9 of the Appendix, we present comparative examples of different methods, highlighting the superior performance of SlideChat. Additionally, Figure 8 in the Appendix showcases examples of SlideChat’s continuous dialogue capabilities, demonstrating its effectiveness in facilitating interactive and comprehensive pathological analysis.
Table 3: VQA performance with different methods. Slide (T) refers to the WSI thumbnail with size of $1024\times 10241024 × 1024$.
MLLMs Input SlideBench-VQA (TCGA) SlideBench-VQA (BCNB) WSI-VQA*
@@ 4 Experiment
Besides, on SlideBench-VQA (BCNB), we compared SlideChat’s zero-shot capabilities with those of different methods. In the zero-shot VQA setting, SlideChat significantly outperforms all other models, achieving the highest overall score of 54.14%. It particularly excels in identifying tumor types, far surpassing other baselines in this task. This performance highlights SlideChat’s generalization capability across a wide range of tasks. When taking patches as inputs, GPT-4o outperforms both MedDr and LLaVA-Med, achieving a score of 41.43%, though it still falls short of SlideChat by 12.71%. Notably, GPT-4o and LLaVA-Med performed very poorly, achieving a score of zero across all tasks when evaluated using slide thumbnails from this testing set. MedDr also shows a notable drop in performance when switching from patch to thumbnail inputs, with its overall score falling from 35.48% to 33.67%. This outcome highlights that, for complex WSIs, relying solely on relatively sufficient visual features is inadequate for effectively supporting a range of tasks. Additionally, in the more fine-grained pathological tasks of the BCNB benchmark, as shown in the Figure 4, SlideChat attains state-of-the-art performance on 5 out of 7 tasks, further demonstrates the effectiveness of SlideChat.
@@
We also curated a subset of closed-set VQA pairs from the public WSI-VQA (Chen et al., 2024a) dataset based on our split test WSI list, referred to as WSI-VQA*, to evaluate the model’s performance. SlideChat demonstrates the highest performance with a score of 60.18. Although MedDr performs well with both patch inputs (54.36%) and slide thumbnail inputs (50.95%), it still falls short compared to SlideChat. GPT-4o struggles significantly, especially with slide thumbnails, scoring only 14.03%, which highlights the limitations of using lower-resolution inputs. SlideChat’s ability to process whole-slide images allows it to leverage both detailed local information and broader context, making it the most effective model for this benchmark. This further emphasizes its superior capability in handling whole-slide data for VQA tasks.
Ablation
We performed ablation experiments from different perspectives as follows: a) Large Language Model Comparison: We compare the performance of several large language models, each with a parameter scale of approximately 7 billion. The evaluated models include Vicuna-7B-v1.5 (Chiang et al., 2023), Phi-3-Mini-4k-Instruct (Abdin et al., 2024), Qwen1.5-7B-Chat (Bai et al., 2023), Llama3-8B-Instruct (AI@Meta, 2024), InternLM2-Chat-7B (Cai et al., 2024), and Qwen2.5-7B-Instruct (Yang et al., 2024). Specifically, we measured their performance on the SlideBench-Caption task using the GPT-score and their accuracy on three VQA (Visual Question Answering) benchmark datasets. As shown in  Table 4, the results demonstrate that SlideChat, powered by the Qwen2.5-7B-Instruct model, achieved the highest performance across all tasks, particularly excelling in the captioning task. These findings underscore the significant potential of developing SlideChat with the Qwen2.5-7B-Instruct model. Utilizing the Qwen2.5 model, we further evaluated models of varying scales and discovered that larger models generally exhibited superior performance, particularly the 7B and 14B parameter models. While these two models showed comparable performance across the three benchmarks, the 14B model surpassed the 7B model by 2.57% on the SlideBench-VQA (TCGA) task. Given computational resource constraints, SlideChat uses the 7B model by default to achieve the best balance between performance and resource efficiency. b) Slide-level encoder effectiveness: We investigate the effectiveness of the slide-level encoder by initially removing it from SlideChat and employing a two-stage training approach. In the first stage, we exclusively trained the projection layers. However, this approach failed to reduce training loss or generate coherent text effectively, likely due to the difficulty of learning the complex visual features of WSIs without the slide-level encoder. With the LLMs frozen and tasked with complex text generation, a simple projection proved insufficient for effectively integrating and aligning visual and textual features. Subsequently, we consider combining data from both stages and training SlideChat without the slide-level encoder by simultaneously updating both the projection layers and the LLM. Under this paradigm, performance on SlideBench-VQA (TCGA) and WSI-VQA (sub), which share the same distribution as the training set, was comparable to the two-stage training configuration with the slide-level encoder. However, a significant decline is observed when evaluating SlideBench-VQA (BCNB), which originates from a different domain; overall performance dropped by over 10% (Table 4), indicating a substantial reduction in the model’s generalization ability. Therefore, we recommend incorporating a slide-level encoder to capture the complex visual features of whole slide, as it is particularly effective for cross-domain alignment and enhances the model’s generalization performance.
@@
Despite SlideChat demonstrating promising results, concerns remained regarding the model’s perception of large pathological slides. To further assess the model’s interpretability, we calculated the correlation between the text output and specific image patches, thereby obtaining patch-level attention scores. By identifying the most significant patches, we gained insights into the precise areas the model focused on during response generation. Highlighting the most relevant regions of the tissue slides not only enhances transparency and bolsters the reliability of the model’s outputs but also assists pathologists by directing attention to critical areas requiring closer scrutiny. Ultimately, such interpretability is essential for fostering trust in AI-assisted diagnostics and enhancing the precision and efficiency of clinical evaluations. As shown in Figure 5, we are pleased to observe that the top five important patches identified by the model closely corresponded with the features described in SlideChat’s output. Our extraction method retrieves attention weights for patch tokens from each generated token, averaging them across layers and heads. We then identify the top five patch tokens with the highest attention weights for further analysis. For example, in Figure 5 (A), the highlighted patches clearly emphasized regions exhibiting an increased nuclear-to-cytoplasmic ratio, hyperchromatic nuclei, and prominent nucleoli. Similarly, in Figure 5 (B), the selected patches demonstrated areas with dense collagen deposition and reduced cellularity, as detailed in the model’s response. This alignment between the highlighted image regions and the textual outputs significantly enhances the model’s interpretability, providing increased confidence that it accurately captures and assesses relevant histopathological features. Such consistency deepens our understanding of the model’s reasoning processes regarding pathological slides and underscores the potential for integrating these AI systems into clinical workflows with greater assurance.
Computational Cost Analysis
@@ 5 Conclusion
In this work, we present SlideChat, the first vision-language assistant capable of understanding gigapixel whole-slide images. Furthermore, we creat SlideInstruction, a largest comprehensive WSI instruction-following dataset to develop SlideChat, as well as SlideBench, a multi-modal benchmark designed to evaluate SlideChat across diverse scenarios. SlideChat demonstrates excellent chat abilities and achieves state-of-the-art performance on 18 tasks.
We bridge the gap between MLLMs and pathology images at the whole-slide level with SlideChat, and believe that it represents a significant advancement towards general pathology and general medical artificial intelligence (GMAI).
Appendix
[…]
Appendix A SlideInstruction and SlideBench
A.1 Data Source
In this section, we present the sources of the constructed SlideInstruction and SlideBench, which are derived from ten TCGA datasets as well as the BCNB challenge dataset. The Table 5 provides a detailed overview of the specific number of WSIs.

### v3

@@ 3.3 Two-stage Training
SlideChat adopts a two-stage training approach (see Fig. 2 B). In the first stage, the primary objective is to align the LLM’s word embeddings with the visual features extracted from whole slide images. This alignment enables the LLM to interpret visual representations from the slide-level encoder, facilitating the effective utilization of the intricate features within the slides. During this stage, SlideChat is trained to generate descriptive captions using 4.2K WSI-caption pairs from SlideInstruction. Specifically, only the slide-level encoder and projection matrix are updated, while the patch-level encoder and LLM weights remain fixed.
@@
In the second stage, we focus on visual question-answering tasks to train the model to accurately respond to domain-specific questions concerning whole slide images. During this phase, the model develops the ability to handle a broad range of multimodal instructions, enabling it to generate answers by effectively integrating both visual and textual information. For example, the model must perform various pathology tasks, such as describing the extent of tumor invasion or assessing the degree of cellular differentiation. To accomplish this, we utilize 176K WSI VQAs from SlideInstruction in the second training stage, allowing the slide encoder, projection layer, and large language model components to be fully trainable to ensure comprehensive adaptability. This approach significantly enhances the model’s capability to handle diverse pathology-related tasks, thereby increasing its effectiveness in real-world clinical and research settings.
4 Experiment
We conducted the following experiments to evaluate three key aspects of SlideChat: (1) its whole slide image captioning capability, which assesses proficiency in generating descriptive captions that accurately summarize the critical pathological features of a WSI; (2) its visual question-answering (VQA) ability across various complex pathological scenarios and its generalizability in zero-shot settings; and (3) SlideChat’s ability to process gigapixel WSIs, capturing both essential global context and intricate details, thereby enhancing its performance compared to patch-level MLLMs. For WSI captioning baselines, we benchmark against MI-Gen [7], a state-of-the-art method specifically designed for this task. Given that existing MLLMs cannot handle the gigapixel scale of whole slide images, we establish baseline comparisons using two approaches: (1) randomly selecting 30 patches from each WSI and inputting them into MLLMs (e.g., GPT-4o [2], LLaVA-Med [17], MedDr [17], Quilt-LLaVA [23]), followed by a majority voting scheme to generate slide-level predictions; and (2) directly inputting a WSI thumbnail, resized to 1024×1024 pixels, referred to as Slide (T), into the MLLMs. For VQA tasks, we further evaluate performance by comparing against random prediction baselines and text-only models, thereby assessing the incremental contribution of visual information. Unless otherwise specified, SlideChat is configured with the patch-level encoder CONCH [20], the slide-level encoder LongNet [12], and utilizes the Qwen2.5-7B-Instruct [33] as LLM.
Table 1: Captioning performance across different methods on SlideBench-Caption.
Methods Input BLEU-1 BLEU-2 BLEU-3 BLEU-4 Rouge-L GPT score
@@ 4 Experiment
Quilt-LLaVA Patch 0.30 0.19 0.11 0.05 0.18 3.87
Quilt-LLaVA Slide (T) 0.23 0.09 0.04 0.01 0.16 1.89
@@
Table 2: VQA performance with different methods.
Methods Input SlideBench-VQA (TCGA) SlideBench-VQA (BCNB) WSI-VQA*
@@
SlideBench-Caption
We report BLEU, ROUGE, and GPT scores to evaluate caption generation performance in  Tab. 1. For the GPT score, we use GPT-4 to assess the similarity between the generated captions and the ground truth, providing an overall score on a scale of 1 to 10, with higher scores indicating better performance. When utilizing patch-level inputs, GPT-4o generates individual descriptions for each patch, which are subsequently integrated to create the final slide-level caption. However, this approach yields poor performance, as evidenced by a BLEU-1 score of 0.16 and a GPT-score of 1.54. These results suggest that the patch-based method fails to adequately capture the broader context necessary for accurate WSI captioning. When the WSI thumbnail of size 1024×1024 pixels is used as input to GPT-4o, performance decreases further, with a BLEU-1 score of 0.10 and a GPT-score of 1.03. This suggests that while the thumbnail offers a global view of the slide, it may lack the resolution and detail necessary for generating precise and informative captions. In contrast, MI-Gen, a model specifically designed for WSI captioning, demonstrates significantly superior performance across all metrics, achieving a BLEU-1 score of 0.37 and a GPT score of 4.14. Similarly, SlideChat, shows competitive results with a BLEU-1 score of 0.37 and a GPT score of 4.11. These outcomes highlight SlideChat’s ability to effectively integrate both local and global information from the slides and confirm its efficiency in describing whole-slide images.
@@
SlideBench-VQA (TCGA)
We further evaluate SlideChat’s overall performance on the multiple-choice VQA benchmark. The results on SlideBench-VQA (TCGA) in Tab. 2, compare different methods across three categories: microscopy, diagnosis, and clinical, along with an overall performance score. Random selection achieves an overall score of 25.02% accuracy, serving as a baseline for answer distribution but demonstrating poor performance. While GPT-4, relying solely on text input, outperforms random predictions, it continues to struggle with accurately answering questions. When GPT-4o incorporates patch-level inputs, its performance improves markedly, reaching a score of 57.91% and underscoring the crucial role of detailed visual data. However, using a WSI thumbnail results in a lower score of 34.07%, as the reduced detail restricts its ability to deliver precise answers. MedDr performs well, achieving an overall score of 67.70% with patch-level inputs, though this drops slightly to 64.25% when using the slide thumbnail due to the loss of fine visual details. SlideChat outperforms all other methods, attaining a leading overall accuracy of 81.17%, excelling across all categories and significantly surpassing the competition. Even in more fine-grained pathological scenarios, as depicted in the left portion of Fig. 4, SlideChat remains the top-performing model across 13 tasks, particularly in areas such as cytomorphological characteristics, histopathological changes, disease detection, disease classification, and staging and grading, which require the identification of complex pathological visual features. Compared with baselines taking some patches or slide thumbnail as inputs, SlideChat has the capability to analyze a significantly greater number of pathological features with enhanced detail, effectively capturing both localized features and overarching global patterns, allowing SlideChat to provide more accurate and nuanced insights into pathological variations.
@@
In the zero-shot VQA setting, SlideChat significantly outperforms all other models, achieving the highest overall score of 54.14% on SlideBench-VQA (BCNB). It particularly excels in identifying tumor types, far surpassing other baselines in this task. This performance highlights SlideChat’s generalization capability across a wide range of tasks. When taking patches as inputs, GPT-4o outperforms both MedDr and LLaVA-Med, achieving a score of 41.43%, though it still falls short of SlideChat by 12.59%. Notably, GPT-4o and LLaVA-Med performed very poorly, achieving a score of zero across all tasks when evaluated using slide thumbnails from this testing set. MedDr and Quilt-LLaVA both achieve moderate results, with MedDr scoring 33.67% on patches and 35.48% on thumbnails, while Quilt-LLaVA attains 32.19% on patches and 41.55% on thumbnails. This outcome highlights that, for complex WSIs, relying solely on relatively sufficient visual features is inadequate for effectively supporting a range of tasks. In the more fine-grained pathological tasks of the BCNB benchmark, SlideChat attains state-of-the-art performance on 5 out of 7 tasks (Fig. 4), further demonstrating the effectiveness of SlideChat.
@@
We also curated a subset of closed-set VQA pairs from the public WSI-VQA [8] dataset, referred to as WSI-VQA*, which exclusively consists of samples that were not in the training set, to evaluate the model’s performance. SlideChat demonstrates the highest performance with a score of 60.18. Although MedDr performs well with both patch inputs (54.36%) and slide thumbnail inputs (50.95%), it still falls short compared to SlideChat. GPT-4o struggles significantly, especially with slide thumbnails, scoring only 14.03%, which highlights the limitations of using lower-resolution inputs. SlideChat leverages both detailed local information and broader context to process WSIs, making it the most effective model for this benchmark. This further emphasizes its superior capability in handling whole-slide data for VQA tasks.
@@
Ablation
We performed ablation experiments from different perspectives as follows: a) Large Language Model Comparison: Firstly, we compare the performance of several large language models, each with a parameter scale of approximately 7 billion, including Vicuna-7B-v1.5 [10], Phi-3-Mini-4k-Instruct [1], Qwen1.5-7B-Chat [5], Llama3-8B-Instruct [4], InternLM2-Chat-7B [6], and Qwen2.5-7B-Instruct [33]. Specifically, we measured their performance on the SlideBench-Caption using the GPT-score and their accuracy on three VQA datasets. As shown in  Tab. 3, the results demonstrate that SlideChat, powered by the Qwen2.5-7B-Instruct model, achieved the highest performance across all tasks, particularly excelling in the captioning task. These findings underscore the significant potential of developing SlideChat with the Qwen2.5-7B-Instruct model. Utilizing the Qwen2.5 model, we further evaluated models of varying scales and discovered that larger models generally exhibited superior performance, particularly the 7B and 14B parameter models. While these two models showed comparable performance across the three benchmarks, the 14B model surpassed the 7B model by 2.57% on the SlideBench-VQA (TCGA). Given computational resource constraints, SlideChat uses the 7B model by default to achieve the best balance between performance and resource efficiency. b) Slide-level encoder effectiveness: We investigate the effectiveness of the slide-level encoder by initially removing it from SlideChat and employing a two-stage training approach. In the first stage, we exclusively trained the projection layers. However, this approach failed to reduce training loss or generate coherent text effectively, likely due to the difficulty of learning the complex visual features of WSIs without the slide-level encoder. When using a frozen LLM for complex text generation, a simple projection proved inadequate for effectively integrating and aligning visual and textual features. Subsequently, we consider combining data from both stages and training SlideChat without the slide-level encoder by simultaneously updating both the projection layers and the LLM. Under this paradigm, performance on SlideBench-VQA (TCGA) and WSI-VQA*, which share the same distribution as the training set, was comparable to the two-stage training configuration with the slide-level encoder. However, a significant decline is observed when evaluating SlideBench-VQA (BCNB), which originates from a different domain; overall performance dropped by over 10% (Tab. 3), indicating a substantial reduction in the model’s generalization ability. Therefore, we recommend incorporating a slide-level encoder to capture the complex visual features of whole slide, as it is particularly effective for cross-domain alignment and enhances the model’s generalization performance.
@@
Despite SlideChat demonstrating promising results, concerns remained regarding the model’s perception of large pathological slides. To further assess the model’s interpretability, we calculated the correlation between the text output and specific image patches, thereby obtaining patch-level attention scores. By identifying the most significant patches, we gained insights into the precise areas the model focused on during response generation. Highlighting the most relevant regions of the tissue slides not only enhances transparency and bolsters the reliability of the model’s outputs but also assists pathologists by directing attention to critical areas requiring closer scrutiny. Ultimately, such interpretability is essential for fostering trust in AI-assisted diagnostics and enhancing the precision and efficiency of clinical evaluations. As shown in  Fig. 5, we are pleased to observe that the top five important patches identified by the model closely corresponded with the features described in SlideChat’s output. Our extraction method retrieves attention weights for patch tokens from each generated token, averaging them across layers and heads. We then identify the top five patch tokens with the highest attention weights for further analysis. For example, in Fig. 5 (A), the highlighted patches clearly emphasized regions exhibiting an increased nuclear-to-cytoplasmic ratio, hyperchromatic nuclei, and prominent nucleoli. Similarly, in Fig. 5 (B), the selected patches demonstrated areas with dense collagen deposition and reduced cellularity, as detailed in the model’s response. This alignment between the highlighted image regions and the textual outputs significantly enhances the model’s interpretability, providing increased confidence that it accurately captures and assesses relevant histopathological features. Such consistency deepens our understanding of the model’s reasoning processes regarding pathological slides and underscores the potential for integrating these AI systems into clinical workflows with greater assurance.
@@ 5 Conclusion
In this work, we present SlideChat, the first vision-language assistant capable of understanding gigapixel whole-slide images. Furthermore, we creat SlideInstruction, a largest comprehensive WSI instruction-following dataset to develop SlideChat, as well as SlideBench, a multi-modal benchmark designed to evaluate SlideChat across diverse scenarios. SlideChat demonstrates excellent chat abilities and achieves state-of-the-art performance on 18 tasks. We bridge the gap between MLLMs and pathology images at the whole-slide level with SlideChat, and we believe this represents an advancement in the field of general medical artificial intelligence (GMAI).
Acknowledgement
The authors acknowledge the funding provided by the National Key Research and Development Program of China (No.2022ZD0160102).
Supplementary Material
6 SlideInstruction and SlideBench
6.1 Data Source
In this section, we present the sources of the constructed SlideInstruction and SlideBench, which are derived from ten TCGA datasets as well as the BCNB challenge dataset. The Tab. 4 provides a detailed overview of the specific number of WSIs.